OpenStudy (anonymous):

how do i rearrange this equation? for j (pi * j^2)/4 = 16

OpenStudy (blurbendy):

start by multiplying by 4 pi * j^2 = 64 divide by pi j^2 = 64 / pi take the square root of both sides j = sqrt[64/pi]

OpenStudy (anonymous):

oh thanks blur.

OpenStudy (blurbendy):

yw

OpenStudy (unklerhaukus):

\[\pm\]